OpenStudy (anonymous):

Your grandmother deposits five hundred dollars into a savings account for you to access when you get older. The account carries a four percent simple annual interest rate. If you want to let the balance grow until it is worth seven hundred dollars, how many years must you wait?

OpenStudy (kymber):

How much interest will you earn per year?

OpenStudy (anonymous):

\[Interest = Principle \cdot Rate ^ {Time} \] We want to find time. Our rate is 100% + 4% = 104% = 1.04 \[$700 = $500 * 1.04 ^ x \] \[\frac{700}{500} = 1.04^x \] \[ln\left(\frac{700}{500}\right) = x \cdot ln(1.04) \] \[x = \frac{ln\left(\frac{700}{500}\right)}{ ln(1.04)} \] \[ x = 8.57 \] So it's going to take 9 years since interest won't collect until a year finishes.